OpenStudy (anonymous):

Find p(5)-p(3) if P(2x)+P(4x)+P(6x)=24x-6

OpenStudy (anonymous):

isnt it P(5)?????

OpenStudy (anonymous):

no the answer is 4 but i dont know the way

OpenStudy (mayankdevnani):

take p as a common

OpenStudy (mayankdevnani):

p(2x+4x+6x)=24x-6 p(12x)=54x-6

OpenStudy (anonymous):

are thats what i am saying ye ya to question me small p hai ya answer mein

OpenStudy (mayankdevnani):

and then find the value of x

OpenStudy (mayankdevnani):

12x-24x=-6 -12x=-6 x=6/12=1/2

OpenStudy (mayankdevnani):

ok @sara123456

OpenStudy (anonymous):

p(5)=(54*5)/12-6 =270/12-6 =135/6-6 =(135-36)/6 =99/6 =33/2 p(3)=(54*3)/12-6 =162/12-6 =81/6-6 =27/2-6 =(27-12)/6 =15/6 =5/2 p(5)-p(3)=33/2-5/2 28/2 14

OpenStudy (anonymous):

i am getting 14

OpenStudy (anonymous):

the answer is 4

OpenStudy (anonymous):

but i will try it once more anyway.thanx

OpenStudy (zarkon):

Show that \(p(x)=2x-2\) satisfies your original equation

OpenStudy (anonymous):

i did it thank you...